# An Efficient Global Optimization Algorithm with Adaptive Estimates of the Local Lipschitz Constants

### v3

@@ 4.2.1 Division and Sampling
During the global search carried out by HALO, we employ the same sampling and division criteria as the DIRECT algorithm [21]. This choice was made due to the robustness and relative simplicity of the DIRECT algorithm, making it easier to understand and implement the code.
@@ 4.2.2 Adaptive Gradients Approximation
In this section we describe how the vector ~$\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}$ from Eq. 5 is computed. After the sampling and division processes, the algorithm collects information about the behavior of the objective function around each partition. To achieve this, a vector is associated with each partition, which describes the intensity of variation of the objective function inside that partition in terms of absolute slopes. This vector adaptively collects the absolute slopes along each coordinate computed based on the points generated so far by the algorithm. We define the nearest neighbors of a partition $\mathcal{D}_{i_{k}}$ as the indices of the nearest points along each coordinate axis.
@@
The nearest neighbors of a partition $\mathcal{D}_{i_{k}}$ are defined as the indices of the nearest points along each coordinate axis. If the partition $\mathcal{D}_{i_{k}}$ is always selected after a certain iteration $\bar{k}$, the nearest neighbors of its centroid $\mathbf{x}_{i_{k}}$ are located at distances ±$\pm\Delta_{i_{k}}$ along each coordinate axis.
@@
|~$|\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}_{n}|=\frac{|f(\mathbf{x}_{i_{k}}^{p_{1}})-f(\mathbf{x}_{i_{k}}^{p_{2}})|}{2\Delta_{i_{k}}},\qquad n=1,\dots,N$ (16)
@@
|~$|\widetilde{\nabla}{f(\mathbf{x}_{i_{k}}^{j})}_{n}|=\begin{cases}\frac{|f(\mathbf{x}_{i_{k}}^{j})-f(\mathbf{x}_{i_{k}})|}{\Delta_{i_{k}}},&\text{if $\mathbf{x}_{i_{k}}\in\mathcal{N}_{i_{k}}^{j}$}.\\ |\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}_{n}|,&\text{otherwise}.\end{cases}$ (17)
For the remaining $N-1$ coordinates, the absolute slopes |~$|\widetilde{\nabla}{f(\mathbf{x}_{i_{k}}^{j})}_{n}|$ are set to the slope of $\mathbf{x}_{i_{k}}$. Although this approach may lead to an incorrect estimation of the gradient associated with the partition $\mathcal{D}_{i_{k}}^{j}$, the effect is negligible as either the partition $\mathcal{D}_{i_{k}}^{j}$ will be selected and its gradient updated or the size of the partition $\mathcal{D}_{i_{k}}^{j}$ is already relatively small.
@@
|~$|\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}_{n}|=\begin{cases}\frac{|f(\mathbf{x}_{i_{k}}^{p_{1}})-f(\mathbf{x}_{i_{k}}^{p_{2}})|}{2\Delta_{i_{k}}},&\text{if$\mathbf{x}_{i_{k}}^{p_{1}},\mathbf{x}_{i_{k}}^{p_{2}}\in\mathcal{N}_{i_{k}}$}.\\ |\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}_{n}|,&\text{otherwise}.\end{cases}$ (18)
@@ 4.3 Coupling Strategy with Local Optimization Methods
$\mathcal{C}_{k}=\{j_{k}:||\mathbf{x}_{i_{k}}-\mathbf{x}_{j_{k}}||\leq r\},\qquad\forall j_{k}\in I_{k},\qquad i_{k}\neq j_{k}$ (20)
@@
$\mathcal{C}_{k}=\mathcal{C}_{k}\cup\{j_{k}:||\mathbf{x}_{p_{k}}-\mathbf{x}_{j_{k}}||\leq r\},\qquad\forall j_{k}\in I_{k},\qquad p_{k}\neq j_{k}$ (23)
@@ 4.4 Convergence and Stopping Criteria
In this paragraph, we will discuss the convergence properties of the HALO algorithm. First, we introduce a property that characterizes the partitions generated during the iterations
@@
The sequence of partitions $\{\mathcal{D}^{i_{k}}\}$ is strictly nested if and only if
@@
According to [13], for the HALO algorithm to converge to a global minimum, the sequence of points generated by the algorithm must be dense in $\mathcal{D}$. We can establish a proposition that shows how the everywhere dense property of HALO is achieved when the largest partitions are always selected
@@
$\lim_{k\to+\infty}||\mathbf{u}_{i_{k}}-\mathbf{l}_{i_{k}}||\leq||\mathbf{u}_{j_{k}}-\mathbf{l}_{j_{k}}||=0$ (28)
with $j_{k}\in\mathcal{I}_{k}^{\text{max}}$, with $\mathcal{I}_{k}^{\text{max}}=\{i_{k}\in I_{k}:||\mathbf{u}_{i_{k}}-\mathbf{l}_{i_{k}}||=\max_{i_{k}\in\mathcal{I}_{k}}||\mathbf{u}_{i_{k}}-\mathbf{l}_{i_{k}}||\}$. ∎
@@
Generally, rather than producing an everywhere dense set we often aim for an algorithm that produces denseness primarily around the global minimizers and not throughout the entire domain $\mathcal{D}$. However, achieving this in practice is very difficult, especially without prior knowledge of the global Lipschitz constant [51, 38]. Hence, the stopping criterion for HALO is mainly based on the maximum number of function evaluations similarly as in [52], which is a commonly used criterion in practice.
@@
Proposition 2 states that the minimum distance $\Delta_{i_{k}}$ from the nearest point to $\mathbf{x}_{i_{k}}$ tends to zero as the iteration $k$ approaches infinity and as the partition $\mathcal{D}_{i_{k}}$ is continuously selected. Thus, as a partition $\mathcal{D}_{i_{k}}$ is selected during the iterations, the information contained in the vector ~$\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}$ becomes more precise, so that we can introduce the following propositions
@@
Given the representative point $\mathbf{x}_{i_{k}}$ which is center of its relative partition $\mathcal{D}_{i_{k}}$. If HALO satisfies has Property 2 and Property 1 then the vector ~$\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}$ associated to the partition $\mathcal{D}_{i_{k}}$ converge to the true gradient at $\mathbf{x}_{i_{k}}$
@@
Proposition 3 further shows that as the selected partition becomes denser the associated vector ~$\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}$ becomes a more accurate approximation of the true gradient $\nabla{f(\mathbf{x}_{i_{k}})}$. As a result, from Proposition 4 the global Lipschitz constant $L$ can be accurately estimated.
@@ 4.5 A Detailed Implementation of HALO
Starting with the counter $k$ set to zero, lines 9 and 10 define the indices of the selected partitions $\mathcal{I}^{\star}_{k}$ and the matrix $\mathbf{S}\in\mathbb{R}^{|\mathcal{I}_{k}|\times N}$, where each entry contains the vector $\mathbf{s}$ defined in Eq. 9. In line 11, a for loop begins, iterating over the indices of each selected partition. This loop is primarily responsible for computing $\Delta_{i^{*}_{k}}$ as described in Section 4.2.1. In line 17, another for loop starts, cycling through each element of the set $\mathcal{P}_{i^{*}_{k}}$ defined in line 14 and Eq. 10. This loop handles the sampling process described in Section 4.2.1. Additionally, we compute the slopes by updating the matrix $\mathbf{G}\in\mathbb{R}^{|\mathcal{I}_{k}|\times N}$, which stores the gradient estimations discussed in Section 4.2.2. The notation $\mathbf{G}_{i^{*}_{k},p}$ indicates that the $p$th coordinate of the gradient relative to the $i^{*}_{k}$th partition is being updated. After this inner for loop, line 28 performs a sorting of the indices of the set $\mathcal{T}_{i^{\star}_{k}}$ defined in Eq. 13, and line 29 calls the function (line 68) responsible for dividing the set $\mathcal{D}_{i^{*}_{k}}$.
@@
Once the outer loop defined in line 11 is completed, based on the information in the matrix $\mathbf{G}$, we invoke the function that computes the values of the local Lipschitz constants, stored in the vector $\mathbf{l}\in\mathbb{R}^{|\mathcal{I}{k}|}$. Lines 31 and 32 define the vectors $\mathbf{v}\in\mathbb{R}^{|\mathcal{I}{k}|}$ and $\mathbf{h}\in\mathbb{R}^{|\mathcal{I}_{k}|}$, which contain half of the distances from the center of each partition $j$ to its vertices and the norm of each gradient estimation (i.e., the norm of each row of $\mathbf{G}$). Line 34 computes the estimation of the local Lipschitz constant in a vectorized form, following Eq. 5.
The next task is to use the local Lipschitz constants collected in the vector $\mathbf{l}$ to compute the lower bounds and select the most promising partitions based on Def. 2. The lower bound values are stored in the vector $\mathbf{r}\in\mathbb{R}^{|\mathcal{I}_{k}|}$ (line 37). The indices $q^{\star}_{1}$, $q^{\star}_{2}$, and $q^{\star}_{3}$ represent the index of the partition with the minimum lower bound (Criterion  1), the partition with the minimum objective function value (Criterion  2), and the partition with the largest diagonal that has the minimum lower bound (Criterion  3). It is possible for one index to satisfy multiple conditions (e.g., $q^{\star}_{1}=q^{\star}_{2}$), so care should be taken to avoid duplicate entries in the set $\mathcal{I}_{k}^{*}$. In lines 42 and 47, we check if a local search has already been performed from the centroid of a partition and if it satisfies Eq. 19. If both conditions are met, the local search is initiated as explained in Section 4.3. Note that the largest hyperrectangle $q^{\star}_{3}$ is always selected (line 52). Finally, the set of selected partitions is returned, and a new loop can begin, iterating over its elements.
1 function Main(max_iter, max_fun_eval, $\mathcal{D}$)
@@ 3       $\mathcal{C}_{k},\mathbf{G},\mathbf{S},\mathbf{f}\leftarrow\emptyset,[[\,\,]],[[\,\,]],[\,\,]$
6                   $\mathbf{l}\leftarrow$ compute_Lipschitz($\mathbf{G}$, $\mathbf{S}$)
7                   $\mathcal{I}^{\star}_{k}\leftarrow$ selection($\mathcal{I}_{k}$, $\mathbf{l}$, $\mathcal{C}_{k}$)
@@ 32                  $\mathcal{I}_{k}\leftarrow$ partitioning($\mathcal{U}_{i^{\star}_{k}},\mathbf{S},\mathbf{V}$)
37       $\mathbf{v}\leftarrow(||\mathbf{S}_{j}||)_{1\leq j\leq|\mathcal{I}_{k}|}$
38       $\mathbf{h}\leftarrow(||\mathbf{G}_{j}||)_{1\leq j\leq|\mathcal{I}_{k}|}$
@@ 39       $\boldsymbol{\alpha}\leftarrow\mathbf{v}/\sqrt{N}$
40       $\mathbf{l}\leftarrow\boldsymbol{\alpha}\odot\mathbf{h}+(1-\boldsymbol{\alpha})\odot\max\{\mathbf{h}\}$
41       return $\mathbf{l}$
@@
44       $\mathbf{r}\leftarrow\mathbf{f}-\mathbf{v}\odot\mathbf{l}$

### v4

@@ 4.2.1 Division and Sampling
During the global search carried out by HALO, we employ the same sampling and division criteria as the DIRECT algorithm [jones1993-JOTA]. This choice was made due to the robustness and relative simplicity of the DIRECT algorithm, making it easier to understand and implement the code.
@@ 4.2.2 Adaptive Gradients Approximation
In this section we describe how the vector $\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}$ from Eq. 4 is computed. After the sampling and division processes, the algorithm collects information about the behavior of the objective function around each partition. To achieve this, a vector is associated with each partition, which describes the intensity of variation of the objective function inside that partition in terms of absolute slopes. This vector adaptively collects the absolute slopes along each coordinate computed based on the points generated so far by the algorithm. We define the nearest neighbors of a partition $\mathcal{D}_{i_{k}}$ as the indices of the nearest points along each coordinate axis.
@@
The nearest neighbors of a partition $\mathcal{D}_{i_{k}}$ are defined as the indices of the nearest points along each coordinate axis. If the partition $\mathcal{D}_{i_{k}}$ is always selected after a certain iteration $\bar{k}$, the nearest neighbors of its centroid $\mathbf{x}_{i_{k}}$ are located at distances $\pm\Delta_{i_{k}}$ along each coordinate axis.
@@
$|\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}_{n}|=\frac{|f(\mathbf{x}_{i_{k}}^{p_{1}})-f(\mathbf{x}_{i_{k}}^{p_{2}})|}{2\Delta_{i_{k}}},\qquad n=1,\dots,N$ (15)
@@
$|\widetilde{\nabla}{f(\mathbf{x}_{i_{k}}^{j})}_{n}|=\begin{cases}\frac{|f(\mathbf{x}_{i_{k}}^{j})-f(\mathbf{x}_{i_{k}})|}{\Delta_{i_{k}}},&\text{if $\mathbf{x}_{i_{k}}\in\mathcal{N}_{i_{k}}^{j}$}.\\ |\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}_{n}|,&\text{otherwise}.\end{cases}$ (16)
For the remaining N1N-1 coordinates, the absolute slopes $|\widetilde{\nabla}{f(\mathbf{x}_{i_{k}}^{j})}_{n}|$ are set to the slope of $\mathbf{x}_{i_{k}}$. Although this approach may lead to an incorrect estimation of the gradient associated with the partition $\mathcal{D}_{i_{k}}^{j}$, the effect is negligible as either the partition $\mathcal{D}_{i_{k}}^{j}$ will be selected and its gradient updated or the size of the partition $\mathcal{D}_{i_{k}}^{j}$ is already relatively small.
@@
$|\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}_{n}|=\begin{cases}\frac{|f(\mathbf{x}_{i_{k}}^{p_{1}})-f(\mathbf{x}_{i_{k}}^{p_{2}})|}{2\Delta_{i_{k}}},&\text{if $\mathbf{x}_{i_{k}}^{p_{1}},\mathbf{x}_{i_{k}}^{p_{2}}\in\mathcal{N}_{i_{k}}$}.\\ |\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}_{n}|,&\text{otherwise}.\end{cases}$ (17)
@@ 4.3 Coupling Strategy with Local Optimization Methods
[…]
@@
[…]
@@ 4.4 Convergence and Stopping Criteria
In this paragraph, we will discuss the convergence properties of the HALO algorithm. Like other DIRECT type approaches, HALO belongs to the class of adaptive partition [pinter1992convergence] and divide the best methods [Sergeyev01011998]. These algorithms typically ensure an everywhere dense type of convergence [torn1989global], which means they can converge to any point in the feasible region, as long as the objective function is continuous. Some algorithms in this class (see Section 2) rely on a reliability parameter for estimating Lipschitz constants. Although a poorly chosen value can impact performance, these methods still enjoy strong theoretical convergence guarantees, even under imperfect parameter tuning [Sergeyev01011998, sergeyev2017deterministic]. HALO avoids the need for this parameter altogether by dynamically estimating local Lipschitz constants at each iteration, together with selecting the largest hyperrectangle, eliminating the need for the user to choose a hyperparameter that affects convergence. This design choice prevents the algorithm from getting stuck in local minima and ensures everywhere dense convergence.
We introduce a property that characterizes the partitions generated during the iterations
@@
The sequence of partitions $\{\mathcal{D}_{i_{k}}\}$ is strictly nested if and only if
@@
We can establish a proposition that shows how the everywhere dense property of HALO is achieved when the largest partitions are always selected
@@
[…]
@@
Generally, rather than producing an everywhere dense set we often aim for an algorithm that produces denseness primarily around the global minimizers and not throughout the entire domain $\mathcal{D}$. However, achieving this in practice is very difficult, especially without prior knowledge of the global Lipschitz constant [di2016direct, sergeyev2015deterministic]. Hence, the stopping criterion for HALO is mainly based on the maximum number of function evaluations similarly as in [liuzzi2010partition], which is a commonly used criterion in practice.
@@
Proposition 2 states that the minimum distance $\Delta_{i_{k}}$ from the nearest point to $\mathbf{x}_{i_{k}}$ tends to zero as the iteration $k$ approaches infinity and as the partition $\mathcal{D}_{i_{k}}$ is continuously selected. Thus, as a partition $\mathcal{D}_{i_{k}}$ is selected during the iterations, the information contained in the vector $\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}$ becomes more precise, so that we can introduce the following propositions
@@
Given the representative point $\mathbf{x}_{i_{k}}$ which is center of its relative partition $\mathcal{D}_{i_{k}}$. If HALO satisfies Property 2 and Property 1 then the vector $\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}$ associated to the partition $\mathcal{D}_{i_{k}}$ converge to the true gradient at $\mathbf{x}_{i_{k}}$
@@
Proposition 3 further shows that as the selected partition becomes denser the associated vector $\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}$ becomes a more accurate approximation of the true gradient $\nabla{f(\mathbf{x}_{i_{k}})}$. As a result, from Proposition 4 the global Lipschitz constant $L$ can be accurately estimated.
@@ 4.5 A Detailed Implementation of HALO
Starting with the counter $k$ set to zero, lines 9 and 10 define the indices of the selected partitions $\mathcal{I}^{\star}_{k}$ and the matrix $\mathbf{S}\in\mathbb{R}^{|\mathcal{I}_{k}|\times N}$, where each entry contains the vector $\mathbf{s}$ defined in Eq. 8. In line 11, a for loop begins, iterating over the indices of each selected partition. This loop is primarily responsible for computing $\Delta_{i^{*}_{k}}$ as described in Section 4.2.1. In line 17, another for loop starts, cycling through each element of the set $\mathcal{P}_{i^{*}_{k}}$ defined at line 14 and Eq. 9. This loop handles the sampling process described in Section 4.2.1. Additionally, we compute the slopes by updating the matrix $\mathbf{G}\in\mathbb{R}^{|\mathcal{I}_{k}|\times N}$, which stores the gradient estimations discussed in Section 4.2.2. The notation $\mathbf{G}_{i^{*}_{k},p}$ indicates that the $p$th coordinate of the gradient relative to the $i^{*}_{k}$th partition is being updated. After this inner for loop, line 28 performs a sorting of the indices of the set $\mathcal{T}_{i^{\star}_{k}}$ defined in Eq. 12, and line 29 calls the function (line 68) responsible for dividing the set $\mathcal{D}_{i^{*}_{k}}$.
@@
Once the outer loop defined in line 11 is completed, based on the information in the matrix $\mathbf{G}$, we invoke the function that computes the values of the local Lipschitz constants, stored in the vector $\boldsymbol{\ell}\in\mathbb{R}^{|\mathcal{I}{k}|}$. Lines 31 and 32 define the vectors $\mathbf{v}\in\mathbb{R}^{|\mathcal{I}{k}|}$ and $\mathbf{h}\in\mathbb{R}^{|\mathcal{I}_{k}|}$, which contain half of the distances from the center of each partition $j$ to its vertices and the norm of each gradient estimation (i.e., the norm of each row of $\mathbf{G}$). Line 34 computes the estimation of the local Lipschitz constant in a vectorized form, following Eq. 4.
[…]
(a) $k=0$
(b) $k=1$
(c) $k=2$
(d) $k=3$
4.5.1 Highlighting the Variable Importance and Problem Interpretability
As discussed in Section 4.5.1, HALO offers more than just an optimization routine; it also provides valuable insights into the importance of each variable in the black-box objective function. After HALO completes its search, we can compute the variable importance using the matrix $\mathbf{G}$, which contains the approximate gradients $\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}$. By averaging these gradients as shown in Equation 33, we identify the directions in which the objective function is most sensitive. In many practical applications, it is not only important to identify the global minimum but also to gain a deeper understanding of the variables that significantly impact the reduction of the objective function. Traditionally, such insights are obtained through sensitivity analysis, which involves computing gradients of the objective function. However, this process can be time-consuming, especially when dealing with computationally expensive simulations.
With HALO, we have an alternative approach to extracting similar information without the need for explicit and additional gradient computations. Once the optimization routine is completed, we can analyze the matrix $\mathbf{G}$ (introduced in Section 4.5), which collects the vectors $\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}$ around each centroid $\mathbf{x}_{i_{k}}$. By examining this matrix, we can gain insights into the directions in which the objective function exhibits greater sensitivity during the iterations of HALO. We can simply determine the variable importance by averaging the rows $\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}$ of the matrix $\mathbf{G}$. This can be expressed as
$\text{Variable Importance}=\frac{1}{|\mathcal{I}_{k}|}\sum_{i}^{|\mathcal{I}_{k}|}\widetilde{\nabla}{f(\mathbf{x}_{i_{k}})}$ (33)
4.5.2 Example of HALO for Univariate Function
5 Numerical Results
In this section, we conduct a comprehensive evaluation of the HALO algorithm compared to other GO algorithms. The evaluation is performed on a diverse set of test functions organized into three different benchmarks. Additionally, we show how to extract insightful information from the objective function using our approach. Finally, we conduct a parametric study on the parameter $\beta$ from Eq. 18 to provide general recommendations to users regarding its value.
5.1 Experimental Setup
5.1.1 Algorithms
The following algorithms are included in our numerical experiments:
  • $\textbf{HALO}_{\text{L-BFGS-B}}$ and $\textbf{HALO}_{\text{SDBOX}}$: These are two versions of HALO that are coupled with different local optimizers. $\text{HALO}_{\text{L-BFGS-B}}$ utilizes a quasi-Newton method for bound-constrained problems (L-BFGS-B) [byrd1995limited], while $\text{HALO}_{\text{SDBOX}}$ uses a derivative-free coordinate descent method with an Armijo-type linesearch for bound-constrained optimization [lucidi2002-COA] (SD-BOX). Both versions of HALO have a single hyperparameter, $\beta$, which is defined in Eq. 18. This parameter determines the minimum size a selected partition must have before a local search can be initiated from its centroid. For our experiments, we fix $\beta$ to $10^{-4}$. In $\text{HALO}_{\text{L-BFGS-B}}$, the gradients used in the Quasi-Newton code are computed using finite difference methods.
[…]
  • DIRECT [jones1993-JOTA]: This is the classical DIRECT algorithm, with a fixed hyperparameter $\epsilon$ set to $10^{-4}|f_{\text{min}}|$, as suggested in [jones1993-JOTA].
  • HPLOR [mockus2017application]: HPLOR is a Pareto-reduced Lipschitzian optimization method similar to the DIRECT algorithm, where only a subset of potentially optimal hyperrectangles is selected. Specifically, it chooses the one with the minimum function value and the one with the largest diagonal. This approach assumes the Lipschitz constant to be either very low or very high. To ensure a fair comparison, we introduced a hybridized version of the algorithm, referred to as Hybrid-PLOR (HPLOR), using the same coupling strategy as HALO and incorporating L-BFGS-B as the local optimizer.
    1function Main(max_iter, max_fun_eval, $\mathcal{D}$)
@@ 3  $\mathcal{C}_{k},\mathbf{G},\mathbf{S},\mathbf{f}\leftarrow\emptyset,[[\,\,]],[[\,\,]],[\,\,]$
[…]
@@ 32       $\mathcal{I}_{k}\leftarrow$ partitioning($\mathcal{U}_{i^{\star}_{k}},\mathbf{S},\mathbf{V}$)
[…]
@@ 39  $\boldsymbol{\alpha}\leftarrow\mathbf{v}/\sqrt{N}$
    40  $\boldsymbol{\ell}\leftarrow\boldsymbol{\alpha}\odot\mathbf{h}+(1-\boldsymbol{\alpha})\odot\max\{\mathbf{h}\}$
[…]
@@
    44  $\mathbf{r}\leftarrow\mathbf{f}-\mathbf{v}\odot\boldsymbol{\ell}$